Andy1234:

Suppose you need to install two heaters, which will be powered by conventional residential power i.e. V=120 V. The first heater with Resistance R1 is to produce P1=450W. The second with resistor R2 is to produce P2=1150 W. If the two heaters are to be wired in series, what should r1 and r2 be?

SemiDefinite:

in series, they share same current (I). \(I = \dfrac{120}{R_1 + R_2} \qquad \square\) \(P_1 = I^2 R_1 = 450 \qquad P_2 = I^2 R_2 = 1150\) \(\implies \dfrac{ R_1 }{450 } = \dfrac{R_2 }{1150 } \qquad \triangle\) https://http.cat/100